# Evaluation of machine-learning models to measure individualized treatment effects from randomized clinical trial data with time-to-event outcomes

### v1

@@ 4.2.2 Individualized treatment effects of adjuvant trastuzumab
[…]
ALASSO $\mathbf{0.677}$ $\mathbf{0.641}$ $0.231$ $0.220$ $\mathbf{0.692}$ $\mathbf{0.659}$ $0.170$ $0.139$
CoxCC $0.590$ $0.507$ $\mathbf{0.039}$ $\mathbf{0.049}$ $0.584$ $0.516$ $0.053$ $0.052$
CoxTime $0.582$ $0.505$ $0.087$ $0.093$ $0.584$ $0.523$ $\mathbf{0.037}$ $\mathbf{0.046}$
IF $0.624$ $0.571$ $0.069$ $0.077$ $0.618$ $0.521$ $0.069$ $0.084$
@@
In Table 6, we further studied the point estimates of expected survival. Patient 3 has a survival time of 8 years and 1 month, with censored follow-up, and received chemotherapy. Patient 4has a survival time of 8 years and 6 months, also with censored follow-up, and received both chemotherapy and trastuzumab. The mean survival probability for all patients is high at 2 years, aligning with the observed survival times of these two patients. The CI are wider for Patient 3 under the control condition. For example, with the CoxCC model, the mean survival probability at 2 years for Patient 3 was $0.822$ with a confidence interval of $[0.64,0.932]$, compared to $0.939$ for Patient 2, with a narrower CI of $[0.862,0.974]$.
Table 6: Mean expected survival probabilities and associated CI at 2 years for 2 patients from the breast cancer cohort on adjuvant trastuzumab.
Patient ALASSO CoxCC CoxTime IF
Patient 3 with control $\underset{[0.823,0.826]}{0.825}$ $\underset{[0.64,0.932]}{0.822}$ $\underset{[0.702,0.912]}{0.824}$ $\underset{[0.760,0.77]}{0.765}$
Patient 4 with treatment $\underset{[0.950,0.951]}{0.951}$ $\underset{[0.862,0.974]}{0.939}$ $\underset{[0.863,0.956]}{0.921}$ $\underset{[0.916,0.920]}{0.918}$
@@ 5 Discussion
In this paper, we compared 2 FNNs methods adapted to survival (CoxCC and CoxTime) and one random forests model (IF) for their ability to predict treatment benefit in the context of time-to-event data, using a linear CoxPH model with an adaptive LASSO penalty as a benchmark method. The comparison was conducted on 2 simulation sets: one introducing nonlinear interaction with treatment and another based on Friedman’s random function generator with non-linear and two-order interactions. Additionally, the models were tested on 2 breast cancer data sets.
@@
The machine learning methods performed well in data generation process 1, especially in terms of discrimination capabilities for IF and calibration for FNNs. For data generation process 2, the performances of the machine learning methods are slightly higher than the reference value in terms of discrimination.
With the first breast cancer data set evaluating the effect of taxane chemotherapy, the machine learning models achieved better prediction results compared to the CoxPH model with an adaptive LASSO penalty, suggesting that the data set has some complex biomarkers-by-treatment interactions. In the second data sets the adaptive lasso model had the best discrimination but the FNN or IF better calibration.
@@
In this paper we used measures specifically adapted to treatment benefit evaluation in the context of time-to-event data. The assessment of treatment benefit prediction models has been an active area of research. Efthimiou et al. (2022) worked on developing measures of discrimination and calibration, extending the work of Klaveren et al. (2017) who introduced the C-for-benefit metric specifically for treatment benefit evaluation. We adapted their approach to time-to-event data by calculating this measure using survival probabilities. Additionally, we modified the RMSE metric to assess treatment benefit for time-to-event data by evaluating the difference between predicted and true treatment benefits, both derived from survival probabilities. Using adapted measures for treatment benefit is of crucial importance, as the standard concordance index was not able to capture the treatment benefit directly. Future work could include associating uncertainty measures with individualized treatment recommendations that could also take model uncertainty into account, for example using ensemble methods.
@@
Our work is centered on the RCT framework, allowing the determination of causal relationships between an intervention and its outcome. However, the methods used in this paper could be adapted beyond RCTs and applied to real-world data cohorts by incorporating propensity scores or other causal approaches. Cui et al. (2023) proposed such an approach using causal forests, a machine learning method used to estimate heterogeneous treatment effects in observational studies.
@@
Further work could focus on the relative contribution of the input variables, as it could enable us to identify specific biomarkers interacting with the treatment, and we could determine the biomarker-treatment score for patients defined as the cross-product between the coefficient of the interactions retained in the training set and their biomarkers. Determining the relative importance of the different input variables is challenging in the case of FNNs. As it is a nonlinear type of model, simply looking at the resulting weights and biases outputted by the model is not possible. Also, FNNs do not have an assumption of non-colinearity of the inputs. There could be informative interactions among groups of variables. To overcome this issue, different methodologies have been defined to assess variable contributions in FNNs that focus on the weights of the model(Garson, 1991) which could be explored. For the IF, we could evaluate the the Effect Importance Measure.
@@ 6 Acknowledgements
The authors acknowledge the National Surgical Adjuvant Breast and Bowel Project (NSABP) investigators of the B-31 trial who submitted data from the original study to dbGaP (dbGaP Study Accession: phs000826.v1.p1) and the NIH data repository. The B-31 trial was supported by: National Cancer Institute, Department of Health and Human Services, Public Health Service, Grants U10-CA-12027, U10-CA-69651, U10-CA-37377, and U10-CA-69974, and by a grant from the Pennsylvania Department of Health.
@@ A Hyperparameter search
[…]
@@ B.3 Detailed results for data generation process 1
[…]
@@
[…]

### v2

@@ 4.2.2 Conditional average treatment effects of adjuvant trastuzumab
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$
ALASSO $\underset{[0.462,0.549]}{0.505}$ $\underset{[0.698,0.756]}{0.729}$ $\underset{[0.062,0.146]}{0.104}$ $\underset{[0.091,0.151]}{0.12}$ $\underset{[0.477,0.533]}{0.503}$ $\underset{[0.693,0.758]}{0.729}$ $\underset{[0.05,0.105]}{0.074}$ $\underset{[0.064,0.111]}{0.086}$
CoxCC $\underset{[0.55,0.625]}{\mathbf{0.588}}$ $\underset{[0.463,0.556]}{0.512}$ $\underset{[0.029,0.106]}{\mathbf{0.065}}$ $\underset{[0.036,0.097]}{0.069}$ $\underset{[0.546,0.607]}{\mathbf{0.581}}$ $\underset{[0.460,0.563]}{0.517}$ $\underset{[0.027,0.084]}{\mathbf{0.058}}$ $\underset{[0.039,0.086]}{0.063}$
CoxTime $\underset{[0.542,0.62]}{0.586}$ $\underset{[0.457,0.556]}{0.504}$ $\underset{[0.033,0.096]}{\mathbf{0.065}}$ $\underset{[0.037,0.094]}{\mathbf{0.064}}$ $\underset{[0.555,0.607]}{0.580}$ $\underset{[0.473,0.564]}{0.512}$ $\underset{[0.021,0.087]}{0.059}$ $\underset{[0.037,0.087]}{\mathbf{0.062}}$
IF $\underset{[0.468,0.535]}{0.503}$ $\underset{[0.715,0.774]}{\mathbf{0.744}}$ $\underset{[0.05,0.141]}{0.088}$ $\underset{[0.08,0.14]}{0.104}$ $\underset{[0.477,0.534]}{0.503}$ $\underset{[0.737,0.787]}{\mathbf{0.762}}$ $\underset{[0.04,0.091]}{0.069}$ $\underset{[0.052,0.099]}{0.076}$
@@
4.2.3 Conditional average treatment effects of avelumab
For the advanced renal cell carcinoma cohort (Table 5), CoxTime achieved the highest concordance at 1 year, indicating stronger discrimination overall, whereas IF obtained the highest $\text{C}_{\text{benefit}}$, suggesting a better ability to capture treatment heterogeneity. In terms of calibration of treatment benefit, CoxCC reached the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, reflecting superior agreement between predicted and observed benefit. By contrast, IF showed higher calibration errors despite its strong discrimination.
Table 5: Bootstrapped evaluation measures at 1 year in the RCT on avelumab. - Matching on predicted benefit.
$t=1$
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$
ALASSO $\underset{[0.341,0.598]}{0.484}$ $\underset{[0.582,0.771]}{0.692}$ $\underset{[0.045,0.289]}{0.179}$ $\underset{[0.079,0.316]}{0.183}$
CoxCC $\underset{[0.417,0.651]}{0.544}$ $\underset{[0.51,0.61]}{0.561}$ $\underset{[0.022,0.14]}{\mathbf{0.061}}$ $\underset{[0.038,0.133]}{\mathbf{0.073}}$
CoxTime $\underset{[0.407,0.664]}{\mathbf{0.549}}$ $\underset{[0.514,0.612]}{0.563}$ $\underset{[0.021,0.144]}{0.068}$ $\underset{[0.036,0.125]}{0.078}$
IF $\underset{[0.39,0.639]}{0.517}$ $\underset{[0.695,0.759]}{\mathbf{0.728}}$ $\underset{[0.158,0.368]}{0.273}$ $\underset{[0.211,0.335]}{0.278}$
@@ 5 Discussion
In this paper, we compared two FNN methods adapted to survival outcomes (CoxCC and CoxTime) and a random forest model (IF) for their ability to predict individualized treatment benefits in the context of time-to-event data, using a linear CoxPH model with an adaptive LASSO penalty as a benchmark method. The comparison was conducted on 2 simulation sets: one introducing nonlinear interaction with treatment and another based on Friedman’s random function generator with nonlinear and two-order interactions. Additionally, the models were evaluated on 3 case studies from oncology clinical trials.
@@
The machine learning methods performed well in data generation process 1, especially in terms of discrimination capabilities for IF and calibration for FNNs. For data generation process 2, the performances of the machine learning methods were better than the linear ALASSO method with regard to the discrimination and calibration metrics, and RMSE.
Across the three clinical trial datasets, the models displayed complementary strengths in predicting treatment benefit. CoxCC and CoxTime generally achieved the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, suggesting more reliable calibration of predicted benefit than IF. IF, on the other hand, tended to yield the highest $\text{C}_{\text{benefit}}$, indicating better identification of patients with larger differential treatment effects, though sometimes at the expense of calibration. ALASSO offered moderate, consistent performance across metrics. Overall, these findings reflect a trade-off: neural network–based models provide well-calibrated benefit estimates, while IF is more sensitive to heterogeneity, and ALASSO may offer a consistent alternative, even if it does not fit nonlinear effects or interactions in these types of cancer trial data sets.
@@
In this paper, we used measures specifically adapted to treatment benefit evaluation in the context of time-to-event data. The assessment of treatment benefit prediction models has been an active area of research. Efthimiou et al. (2022) developed measures of discrimination and calibration, building upon the work of Klaveren et al. (2017), who introduced the C-for-benefit metric specifically for evaluating treatment benefits. We adapted their approach to time-to-event data by calculating this measure using survival probabilities. While C-for-benefit provides a flexible framework for comparing individualized treatment effects, its performance can vary with the matching procedure used and, as noted in recent methodological work (Efthimiou et al., 2025), it does not constitute a proper scoring rule. Additionally, we modified the RMSE metric to assess treatment benefit for time-to-event data by evaluating the difference between predicted and true treatment benefits, both derived from survival probabilities. Using adapted measures for treatment benefit is of crucial importance, as, for example, the standard concordance index is unable to directly capture the treatment benefit. Future work could include associating uncertainty measures with conditional average treatment effects, which could, for example, be evaluated by a bootstrap method as in Roblin et al. (2024).
@@
Weberpals et al. (2025) emphasize the significant potential of leveraging complex causal machine learning models to refine individualized treatment effect estimation. Causal machine learning provides a promising framework for enhancing precision medicine by predicting patient-specific cancer treatment responses using rich, multimodal data. Although clinical applications of causal machine learning are still emerging, its capability to integrate high-dimensional data and provide robust causal inference shows great promise for improving personalized treatment choices. Our work is indeed centered on the RCT framework, which allows for the determination of causal relationships between an intervention and its outcome. However, the methods used in this paper could be adapted beyond RCTs and applied to real-world data cohorts by incorporating propensity scores or other causal approaches. Cui et al. (2023) proposed such an approach using causal forests, a machine learning method used to estimate heterogeneous treatment effects in observational studies.
@@
Further work could focus on the relative contribution of the covariates, as it could enable us to identify specific biomarkers that predict larger treatment effects that could be further investigated in basic or clinical research project. Determining the relative importance of the different input variables is challenging in the case of FNNs. As it is a nonlinear type of model, simply examining the resulting weights and biases output by the model is not possible. Additionally, FNNs do not require the assumption of non-collinearity among the inputs. There could be informative interactions among groups of variables. To overcome this issue, different methodologies have been defined to assess variable contributions in FNNs that focus on the weights of the model (Garson, 1991) which could be explored. For the IF, we could evaluate the Effect Importance Measure. Feature importance could also be evaluated in both types of models using methods based on resampling and permutation Fisher et al. (2019).
6 Conclusion
Neural networks provide well-calibrated treatment benefit predictions, while random forests provide superior discrimination for identifying patients with heterogeneous responses. These complementary strengths reveal an inherent modeling trade-off. Our adapted metrics for survival benefits uncover performance differences that standard indices miss. These findings confirm causal machine learning’s promise for precision medicine, with methods applicable beyond clinical trials. Future work should quantify prediction uncertainty and identify key treatment-response biomarkers.
@@ 7 Acknowledgements
Elvire Roblin acknowledges financial support by Foundation Philanthropia Lombard-Odier. The authors also acknowledge the National Surgical Adjuvant Breast and Bowel Project (NSABP) investigators of the B-31 trial who submitted data from the original study to dbGaP (dbGaP Study Accession: phs000826.v1.p1) and the NIH data repository. The B-31 trial was supported by: National Cancer Institute, Department of Health and Human Services, Public Health Service, Grants U10-CA-12027, U10-CA-69651, U10-CA-37377, and U10-CA-69974, and by a grant from the Pennsylvania Department of Health.
@@ A Hyperparameter search
Optimization algorithm $\{\text{Adam, AdamAMSGRAD, RMSProp, SGDWR}\}$
@@ B.3 Detailed results for data generation process 1
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE
@@
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE